# OC3D: Weakly Supervised Outdoor 3D Object Detection with Only Coarse Click Annotation

### v1

@@ Abstract
LiDAR-based outdoor 3D object detection has received widespread attention. However, training 3D detectors from the LiDAR point cloud typically relies on expensive bounding box annotations. This paper presents OC3D, an innovative weakly supervised method requiring only coarse clicks on the bird’s eye view of the 3D point cloud. A key challenge here is the absence of complete geometric descriptions of the target objects from such simple click annotations. To address this problem, our proposed OC3D adopts a two-stage strategy. In the first stage, we initially design a novel dynamic and static classification strategy and then propose the Click2Box and Click2Mask modules to generate box-level and mask-level pseudo-labels for static and dynamic instances, respectively. In the second stage, we design a Mask2Box module, leveraging the learning capabilities of neural networks to update mask-level pseudo-labels, which contain less information, to box-level pseudo-labels. Experimental results on the widely used KITTI and nuScenes datasets demonstrate that our OC3D with only coarse clicks achieves state-of-the-art performance compared to weakly-supervised 3D detection methods. Combining OC3D with a missing click mining strategy, we propose a OC3D++ pipeline, which requires only 0.2% annotation cost in the KITTI dataset to achieve performance comparable to fully supervised methods.
@@ Introduction
Figure 1: (a) Fully-supervised annotation. (b) In contrast to fully-annotated, coarse-click merely needs coarse clicks on the BEV of objects. (c) Performance comparison of different strategies. The blue bars: inferring box pseudo-labels for all clicks by our Click2Box module. The orange bars: discarding pseudo-labels of dynamic instances after motion classification. The yellow bars: processing dynamic instances by the Click2Mask module. The green bars: the final performance of our PC3D.
@@
To provide a faster, albeit less precise, method of human supervision, WS3D (Meng et al. 2021) and ViT-WSS3D (Zhang et al. 2023) propose a mixed supervision strategy that replaces some box annotations with the center-click annotations. In this approach, annotators click the center of objects on the Bird’s Eye View (BEV) to generate center-level labels, reducing labeling time per instance to approximately 2.5 seconds - 50 times faster than traditional bounding box labeling.
@@
In this paper, we introduce a novel pure click-supervised approach for 3D object detection (OC3D), which employs temporal cues to distinguish between static and moving instances and progressively recover their box supervision from coarse clicks (Fig. 1 (b)) for detector training. This strategy offers greater flexibility than center-clicked approaches by accommodating inaccurate clicks through shift tolerance, which can be corrected via multiple-frame consistency. Our framework comprises three key designs: (1) To efficiently classify the motion state of provided clicks, we analyze the point density at the clicked location from a statistical perspective, noting that vary point density over a short period for moving objects while stable density for static instances. (2) For static objects, we implement a Click2Box strategy that aggregates the neighboring points across multiple frames to reconstruct the object’s structure and lift click-level supervision to a precise 3D box. (3) For moving objects, which are more challenging to regress, we generate mask-level pseudo labels (Click2Mask) to assign the clustered points with labels, enabling models to learn to predict the object’s location. To compensate for the lack of supervision in predicting shapes and scales for moving objects, we leverage their geometric similarity as a bridge, further refining the 3D box through a Mask2Box process. To address extreme cases where only one click is provided per scene (sparsely supervised setting), we propose a variant OC3D++, which identifies missed click instances and facilitates training with only 0.2% of the annotations required for full labeling.
We evaluated OC3D and OC3D++ on the widely adopted KITTI and nuScenes datasets. Remarkably, OC3D achieves competitive performance with weakly-supervised baselines that rely on accurate box annotations. Furthermore, when the number of clicks is reduced to one per scene, OC3D++ still maintains robust performance, demonstrating the resilience and effectiveness of our approach under the sparely supervised condition. In summary, our contributions are:
  • We propose the first method of only click annotated 3D object detection from point cloud (OC3D), which solely relies on coarse clicks on the Bird’s Eye View (BEV). This approach dramatically reduces the annotation cost of 3D object detection tasks.
@@
In recent years, fully-supervised 3D object detection has been widely studied. The early methods (Lang et al. 2019; Yan, Mao, and Li 2018; Yin, Zhou, and Krähenbühl 2021) utilized an end-to-end one-stage object detection strategy, predicting detection boxes directly from point clouds. Subsequently, the two-stage methods (Shi, Wang, and Li 2019; Deng et al. 2021; Shi et al. 2023; Wu et al. 2022b; Chen et al. 2023; Wang et al. 2024) introduced an additional proposal box refinement stage, which improves detection performance by refining regions of interest. Meanwhile, multi-stage methods (Wu et al. 2022a; Cai et al. 2022) have begun to show their potential and prominence. These methods progressively refine proposal boxes using a cascaded optimization strategy, leading to more precise prediction results. Despite achieving excellent detection results, all these methods require fully annotated box-level annotations, the generation of which is time-consuming and labor-intensive.
Figure 2: The pipeline of the proposed PC3D. (a) Initially, a novel motion state classification strategy is introduced, followed by the generation of mask-level pseudo-label $A_{m}$ and box-level pseudo-label $A_{b}$, utilizing the Click2Mask and Click2Box modules, respectively. (b) With the mixed pseudo-labels generated by stage (a), train the detector and then update the mask-level supervision to box-level based on high-confidence predictions of the trained detector.
@@
$\mathfrak{L}=\frac{1}{A_{b}}\sum_{A_{b}}\mathfrak{L}_{\operatorname{reg}}+\frac{1}{A_{b}+A_{m}}\sum_{A_{b}+\left|A_{m}\right|}\mathfrak{L}_{\operatorname{cls}}+\lambda\frac{1}{A_{m}}\sum_{\left|A_{m}\right|}\mathfrak{L}_{\operatorname{pos}}.$ (5)
@@
Figure 3: Demonstration of the mask-level pseudo-label upgrading strategy.
@@
Recently, weakly supervised 3D object detection methods(Xia et al. 2023b, 2024) with sparse bounding boxes have made great progress. The sparse annotation approach can greatly reduce the annotation cost. Motivated by this, we further explore the weakly supervised 3D object detection scheme based on sparse clicks. Unlike conventional click annotations, under sparse clicking, only one instance is randomly selected for annotation in each scene, presenting a greater challenge to the 3D object detector. To address the problem of missing annotation information, we propose a OC3D-based missing-click instances mining method, OC3D++ (for specific design details, see the supplementary materials). The experimental results of OC3D++ demonstrate that our OC3D can be effectively combined with missing-click instances mining methods.
@@
We also conducted validation experiments on the nuScenes dataset. To ensure a fair comparison, we follow the previous methods (Xia et al. 2023b, 2024) to selected the CenterPoint (Yin, Zhou, and Krähenbühl 2021) as the base detector. In Tab. 2, where OC3D achieves performance comparable to the fully supervised baseline method. This confirms that our SC3D still demonstrates outstanding performance on highly challenging multi-class tasks. In the experimental with sparse coarse click annotations, the mAP of OC3D++ decreased. The reason is the higher number of instances in the NuScenes, which poses a greater challenge for sparse settings.
@@
Figure 4: Prediction results and the quality of pseudo-labels across various iterative rounds. (a) and (b) represent the 3D and BEV detection results for car at IOU thresholds of 0.5 and 0.7, respectively. (c) compares the box-level pseudo-labels with the ground truth across each iterative round.
@@
Tab. 4 demonstrates the robustness of the coarse click. To rigorously verify the impact of perturbation factor, we conduct experiments on the more challenging sparse clicking setup. We applied various disturbance factors $\delta=0.25,0.5,0.7,1.0$, defining the disturbance range as $(\delta\times w,\delta\times l)$ , where $w$ and $l$ are the size of the object. Even with a large disturbance range ($\delta=1.0$), our OC3D++ still achieves satisfactory performance. In Tab.5, we discuss the hyper-parameter $\lambda$ in Equ.5 without the Mask2Box modules. Lower weights for ambiguous location predictions yield better results. This result indicates that for ambiguous location supervision, providing a single hint is sufficient to achieve optimal detection performance, and excessive weighting introduces noise to the detector.

### v2

@@ Abstract
LiDAR-based outdoor 3D object detection has received widespread attention. However, training 3D detectors from the LiDAR point cloud typically relies on expensive bounding box annotations. This paper presents OC3D, an innovative weakly supervised method requiring only coarse clicks on the bird’s eye view of the 3D point cloud. A key challenge here is the absence of complete geometric descriptions of the target objects from such simple click annotations. To address this problem, our proposed OC3D adopts a two-stage strategy. In the first stage, we initially design a novel dynamic and static classification strategy and then propose the Click2Box and Click2Mask modules to generate box-level and mask-level pseudo-labels for static and dynamic instances, respectively. In the second stage, we design a Mask2Box module, leveraging the learning capabilities of neural networks to update mask-level pseudo-labels, which contain less information, to box-level pseudo-labels. Experimental results on the widely used KITTI and nuScenes datasets demonstrate that our OC3D with only coarse clicks achieves state-of-the-art performance compared to weakly-supervised 3D detection methods. Combining OC3D with a missing click mining strategy, we propose an OC3D++ pipeline, which requires only 0.2% annotation cost in the KITTI dataset to achieve performance comparable to fully supervised methods. The code will be made publicly available.
@@ Introduction
Figure 1: (a) 6DoF box annotation; (b) Unlike costly box annotations, coarse-click alternatives require only quick clicks on objects in the 2D BEV plane, yet offer limited supervision; (c) An empirical study on maximizing the utility of weak click labels reveals significant performance gains when separating static and dynamic objects (yellow bars) compared to directly estimating 3D bounding boxes from clicks (blue bars) or discarding pseudo labels for dynamic objects (orange bars). The proposed OC3D (green bars) further refines label quality via Mask2Box conversion, substantially enhancing detection performance by 17%.
@@
To provide a faster, albeit less precise, method of human supervision, WS3D (Meng et al. 2021) and ViT-WSS3D (Zhang et al. 2023) propose a mixed supervision strategy that replaces some box annotations with the center-click annotations. In this approach, annotators click the center of objects on the Bird’s Eye View (BEV) to generate center-level labels, reducing labeling time per instance to approximately 2.5 seconds, which is 50 times faster than traditional bounding box labeling.
@@
In this paper, we introduce a novel pure click-supervised approach for 3D object detection (OC3D), which employs temporal cues to distinguish between static and moving instances and progressively recover their box supervision from coarse clicks (Fig. 1 (b)) for detector training. This strategy offers greater flexibility than center-clicked approaches by accommodating inaccurate clicks through shift tolerance, which can be corrected via multiple-frame consistency. Our framework comprises three key designs: (1) To efficiently classify the motion state of provided clicks, we analyze the point density at the clicked location from a statistical perspective, noting that vary point density over a short period for moving objects while stable density for static instances. (2) For static objects, we implement a Click2Box strategy that aggregates the neighboring points across multiple frames to reconstruct the object’s structure and lift click-level supervision to a precise 3D box. (3) For moving objects, which are more challenging to regress, we generate mask-level pseudo labels (Click2Mask) to assign the clustered points with labels, enabling models to learn to predict the object’s location. To compensate for the lack of supervision in predicting shapes and scales for moving objects, we leverage their geometric similarity as a bridge, further refining the 3D box through a Mask2Box process. To address extreme cases where only one click is provided per scene (sparsely supervised setting), we propose a variant OC3D++, which identifies missed click instances and facilitates training with only 0.2% of the annotations required for full labeling. We evaluated OC3D and OC3D++ on the widely adopted KITTI and nuScenes datasets. Remarkably, OC3D achieves competitive performance with weakly-supervised baselines that rely on accurate box annotations. Furthermore, when the number of clicks is reduced to one per scene, OC3D++ still maintains robust performance, demonstrating the resilience and effectiveness of our approach under sparely supervised conditions.
In summary, our contributions are:
  • We propose the first method of only click annotated 3D object detection from point cloud (OC3D), which solely relies on coarse clicks on the BEV maps. This approach dramatically reduces the annotation cost of 3D object detection tasks to 1% $\sim$ 0.2%.
@@
In recent years, fully-supervised 3D object detection has been widely studied. The early methods (Lang et al. 2019; Yan, Mao, and Li 2018; Yin, Zhou, and Krähenbühl 2021) utilized an end-to-end one-stage object detection strategy, predicting detection boxes directly from point clouds. Subsequently, the two-stage methods (Shi, Wang, and Li 2019; Deng et al. 2021; Shi et al. 2023; Wu et al. 2022b; Chen et al. 2023; Wang et al. 2024) introduced an additional proposal box refinement stage, which improves detection performance by refining regions of interest. Meanwhile, multi-stage methods (Wu et al. 2022a; Cai et al. 2022) progressively refined proposals using a cascaded optimization strategy, leading to more precise predictions. Despite achieving excellent performance, all these methods require costly box annotations, the generation of which is time-consuming and labor-intensive.
Figure 2: The pipeline of the proposed OC3D. (a) Initially, a novel motion state classification strategy is introduced, followed by the generation of mask-level pseudo-label $\mathbf{A}_{m}$ and box-level pseudo-label $\mathbf{A}_{b}$, utilizing the Click2Mask and Click2Box modules, respectively. (b) With the mixed pseudo-labels generated by stage (a), train the detector and then update the mask-level supervision to box-level based on high-confidence predictions of the trained detector.
@@
$\mathfrak{L}=\frac{1}{A_{b}}\sum_{A_{b}}\mathfrak{L}_{\operatorname{reg}}+\frac{1}{A_{b}+A_{m}}\sum_{A_{b}+A_{m}}\mathfrak{L}_{\operatorname{cls}}+\lambda\frac{1}{A_{m}}\sum_{A_{m}}\mathfrak{L}_{\operatorname{pos}}.$ (5)
@@
Figure 3: An illustration of the upgrading strategy for mask-level pseudo labels.
@@
Recently, weakly supervised 3D object detection methods (Xia et al. 2023b, 2024) with sparse bounding boxes have made great progress. The sparse annotation approach can greatly reduce the annotation cost. Motivated by this, we further explore the weakly supervised 3D object detection scheme based on sparse clicks. Unlike conventional click annotations, under sparse clicking, only one instance is randomly selected for annotation in each scene, presenting a greater challenge to the 3D object detector. To address the problem of missing annotation information, we propose an OC3D-based missing-click instances mining method, OC3D++ (for specific design details, see the supplementary materials). The experimental results of OC3D++ demonstrate that our OC3D can be effectively combined with missing-click instances mining methods.
@@
We also conducted validation experiments on the nuScenes dataset. To ensure a fair comparison, we follow the previous methods (Xia et al. 2023b, 2024) to select the CenterPoint (Yin, Zhou, and Krähenbühl 2021) as the base detector. In Tab. 2, where OC3D achieves performance comparable to the fully supervised baseline method. This confirms that our SC3D still demonstrates outstanding performance on highly challenging multi-class tasks. In the experimental with sparse coarse click annotations, the mAP of OC3D++ decreased. The reason is the higher number of instances in the nuScenes, which poses a greater challenge for sparse settings.
@@
Figure 4: Prediction results and the quality of pseudo-labels across various iterative rounds. (a) and (b) represent the 3D and BEV detection results for cars at IOU thresholds of 0.5 and 0.7, respectively; (c) compare the box-level pseudo-labels with the ground truth across each iterative round.
@@
Tab. 4 demonstrates the robustness of the coarse click. To rigorously verify the impact of the perturbation factor, we conduct experiments on the more challenging sparse clicking setup. We applied various disturbance factors $\delta=0.25,0.5,0.7,1.0$, defining the disturbance range as $(\delta\times w,\delta\times l)$ , where $w$ and $l$ are the size of the object. Even with a large disturbance range ($\delta=1.0$), our OC3D++ still achieves satisfactory performance. In Tab.5, we discuss the hyper-parameter $\lambda$ in Equ.(5) without the Mask2Box modules. Lower weights for ambiguous location predictions yield better results. This result indicates that for ambiguous location supervision, providing a single hint is sufficient to achieve optimal detection performance, and excessive weighting introduces noise to the detector.